Recent Log Entries

The following cavity histories have been updated: 1L09-4

Log Number: 4131281Logbook(s): Tags: Body: 

The following cavity histories have been updated:
1L09-4
Username: kerns
Date/Time: 02-14-23_21:46:37
RF Status: Good
Tuner Status: Good
Comment: Replaced filament bd. twice and cleaned light pipe and reseated AIB chips. Set filaments down to 5.8v fm 6.8v.
Special Instructions:

1L09-4 Persistent Filament Volts High Drop

Log Number: 4131279Logbook(s): Entry Makers: kernsBody: 

Replaced filament board twice with fil. high trip both times Swapped 3 and 4 board and 4 tripped again. Did a reseat of chips on the AIB and cleaned light pipe ends and filament count kept getting higher until about the eighth time it finally made it to 7.0 KFVM with a setpoint of 6.7v. Put into HV and RF until cav. 4 dropped on high filaments again 12 minutes later. Lowered filament volt set to 5.8v for a KFVM of 6.0 volts. KCCU went from 1.27a to 1.19a but that is more than enough current for the power. Filaments should not be raised to 7.0 volts or else the drops.

Restart iocsoftig2NL13 NDX CAS

Log Number: 4131238Logbook(s): Body: 

Control noticed soft iocs hosted on opsbat0 went unresolved. Theo Larrieu identified iocsoftig2NL13 has a bad player, maybe. So we killed the soft ioc and let the ForeverFile daemon restart it. The controller may need to
be power-cycled in the AM if the HV readback and Temp readback don't pick up.

Status Update

Log Number: 4131217Logbook(s): Entry Makers: bfreemanBody: 

IL09, 1L12 still down. EES (Farrish and Kerns) still working on them.

We are also still having an issue with a soft ioc host opsbat0. Many of the PVs are not resolving on different soft iocs. Sue Witherspoon is working with Sys Admin to get this strightened out.

The following cavity histories have been updated: 1L09-4

Log Number: 4131216Logbook(s): Tags: Body: 

The following cavity histories have been updated:
1L09-4
Username: kerns
Date/Time: 02-14-23_19:44:49
RF Status: Bad
Tuner Status: Good
Comment: Filaments faulting high at 6.9v at about 3v filament count up. Board changed twice.
Special Instructions:

The following cavity histories have been updated: Zone 1L11

Log Number: 4131215Logbook(s): Tags: Body: 

The following cavity histories have been updated:
Zone 1L11
Username: kerns
Date/Time: 02-14-23_19:43:13
RF Status: Good
Comment: CPS fan not on while in HV, dropped both fans and transformer heated. Recovered 2 hours later.
Special Instructions:

Special Instructions:

iocsoftnl22srfvac restored

Log Number: 4131205Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2023-02-14 19:10:38 The output from the myRestore execution was as follows: iocsoftnl22srfvac systemEnum@2023-02-14 19:10 2(2) restored srfvacMB_v3@2023-02-14 19:10 35(35) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsoftnl22srfvac (success) Procedure return status: Success

CREW CHIEF SIGN IN

Log Number: 4131173Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 02-14-23
Time : 18:22:20
Day : Tuesday
Shift : Swing

Crew Chief : Freeman
Program Deputy: Letta
Team : -
Arm : Anthony
Designated SSO: Howard
LERF Operator:
Operators : Anthony, Schaumloffel, Howard

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

COMMENT ON LAM3C Earth Leakage Fault Investigation

Log Number: 4131149Logbook(s): Body: 

Jason, Mark and I investigated the ground fault trip this afternoon and we found that the internal ground fault circuit breaker was tripped. All box power supplies have two mode of measuring and initiating the fault. One is done by an actual ground fault sensor and second one is using an inline circuit breaker (1A). The ground fault trip on LAM3C sensor is set to be around 100mA (historical number). This afternoon trip had also faulted the 1A circuit breaker, which usually means it was a hard fault. A Hard faults are usually easier to recreate as they will keep tripping the circuit breaker.

iocsoftnl10rf restored

Log Number: 4131142Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2023-02-14 06:32:00 The output from the myRestore execution was as follows: iocsoftnl10rf <NOTE> rf3int: Provided macro *IOC* is never used <NOTE> rf3hpa: Provided macro *IOC* is never used <NOTE> rf3stp: Provided macro *IOC* is never used <NOTE> rf3fcc1: Provided macro *IOC* is never used <NOTE> rf3fcc2: Provided macro *IOC* is never used <NOTE> rf3fcc3: Provided macro *IOC* is never used <NOTE> rf3fcc4: Provided macro *IOC* is never used <NOTE> rf3fcc5: Provided macro

iocnl13diag restored

Log Number: 4131139Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2023-02-14 14:15:51 The output from the myRestore execution was as follows: iocnl13diag system@2023-02-14 14:15 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocnl13diag (success) Procedure return status: Success

Pages